# Locally rigid $\infty$-categories

### v1

@@ 4.3.1 Over $\mathbf{Sp}$
Since $\mathcal{V}_{i}$ is rigid, this implies that it is trace-class, and thus the whole of $(x_{i}^{r})_{i}\to(x_{i}^{s})_{i}$ is trace-class in $\prod_{i}\mathcal{V}_{i}$. By the usual trick, it implies that $\mathrm{colim}_{r\in\mathbb{Q}_{\geq 0}}y((x_{i}^{r})_{i})$ is also a $\mathbb{Q}_{\geq 0}$-indexed colimit of things in $\prod_{i}^{\mathrm{dbl}}\mathcal{V}_{i}$ along trace-class maps, and thus it follows that $\prod_{i}^{\mathrm{dbl}}\mathcal{V}_{i}$ is rigid by 4.57.
Finite limits: The empty limit is the terminal object, i.e. 00, and this is clearly rigid. So we are left with pullbacks: let $\mathcal{V}_{0}\xrightarrow{p_{0}}\mathcal{V}_{01}\xleftarrow{p_{1}}\mathcal{V}_{1}$ be a span in $\mathrm{CAlg}^{\mathrm{rig}}$, and $\mathcal{V}$ be its pullback in $\mathrm{CAlg}((\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})^{\mathrm{dbl}})$, $\mathcal{V}^{\prime}$ its pullback in $\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})$, and $\overset{\rightarrow}{\mathcal{V}}$ the doubly lax pullback as in 4.89.
@@
Furthermore, we have observed in 4.91 that the bottom left vertical map is an isomorphism, and it is clear from 4.48 that the top left vertical map also is. Thus it suffices to prove that the image of $\hat{y}:\mathcal{V}\to\mathrm{Ind}(\mathcal{V})\to\mathrm{Ind}(\overset{\rightarrow}{\mathcal{V}})$ is contained in the image of $\mathrm{Rig}(\overset{\rightarrow}{\mathcal{V}})$.
However we already know that the image of $\mathcal{V}$ in $\mathrm{Ind}(\mathcal{V}^{\prime})$ is generated under colimits by $\mathbb{Q}$-indexed colimits of maps between representables whose projections to each $\mathcal{V}_{i}$ are trace-class. Since each map in $\mathbb{Q}$ is a composite of two maps, it follows from 4.89 that each of the maps in question is trace-class in $\overset{\rightarrow}{\mathcal{V}}$, and hence, that the relevant $\mathbb{Q}$-colimits are in $\mathrm{Rig}(\overset{\rightarrow}{\mathcal{V}})$, as was to be shown. ∎
@@ 4.3.2 In general
We will prove in Section 5 that the category of rigid $\mathcal{V}$-algebras is presentable, and explain how to obtain from this a different proof of the existence of rigidifications. However, it is still not explicit enough to obtain 4.63 - the goal of this subsection is to provide an actual construction of the rigidification for which we have enough control to prove this result. This construction is motivated by the construction over $\mathbf{Sp}$, and just as in this case, we first need an “intrinsic” description of rigid categories, obtained as 4.94.
Our goal is ultimately to prove that for a locally rigid $\mathcal{V}$-algebra $\mathcal{W}$, the canonical functor $\mathrm{Rig}_{\mathcal{V}}(\mathcal{W})\to\mathcal{W}$ admits a fully faithful left adjoint. We have already noted that $\mathrm{Rig}_{\mathcal{V}}(\mathcal{W})\to\mathcal{W}$ factors in a fully faithful way through $\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})$, so we split the assignment in two halves: a- for a $\tilde{\mathcal{W}}$ with $\mathcal{V}$-atomic unit (to be thought of as $\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})$, so in particular, not assumed locally rigid), fully describe $\mathrm{Rig}_{\mathcal{V}}(\tilde{\mathcal{W}})$ as a full subcategory of $\tilde{\mathcal{W}}$; b- Use this description to prove that for $\tilde{\mathcal{W}}:=\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})$, $\hat{y}:\mathcal{W}\to\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})$ factors through $\mathrm{Rig}_{\mathcal{V}}(\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa}))=\mathrm{Rig}_{\mathcal{V}}(\mathcal{W})$. The second step will be rather easy once the first is established, so we now focus on the first.
We already noted in 4.71 that if $\mathbf{1}_{\mathcal{W}}$ is atomic, then $\mathrm{Rig}_{\mathcal{V}}(\mathcal{W})\to\mathcal{W}$ is fully faithful, so our goal is to describe the essential image.
@@
Suppose $\mathcal{W}$ is a commutative $\mathcal{V}$-algebra with $\mathcal{V}$-atomic unit. $\mathcal{W}$ is rigid if and only if it satisfies an analogue of 2.25 with trace-class presentability, that is, if and only if:
@@
As in the case of 2.25, it is worth starting with a preparatory lemma:
@@
4.95 shows that if $f:\mathcal{M}\to\mathcal{N}$ is a lax $\mathcal{W}$-linear functor which is strong $\mathcal{V}$-linear and colimit-preserving, then any of the assumptions 1.-4. or their “generated by” analogues imply that $f$ is strong $\mathcal{W}$-linear.
@@
First we observe that there is a map $\mathrm{trcl}_{\mathcal{W}}^{\mathcal{V}}\to{\mathrm{at}}_{\mathcal{W}}^{\mathcal{V}}$ since $\mathbf{1}_{\mathcal{W}}$ is atomic. Therefore, our assumptions imply the assumptions of 2.25, so that $\mathcal{W}$ is dualizable over $\mathcal{V}$. We may therefore apply 2.31 to check whether this is a $\mathcal{V}$-internal left adjoints.
Now we note that tensor products of trace-class maps are trace-class - a refinement of this is the existence of a transformation $\mathrm{trcl}_{\mathcal{W}}^{\mathcal{V}}(w_{0},w_{1})\otimes\mathrm{trcl}_{\mathcal{W}}^{\mathcal{V}}(w_{0}^{\prime},w_{1}^{\prime})\to\mathrm{trcl}_{\mathcal{W}}^{\mathcal{V}}(w_{0}\otimes w_{0}^{\prime},w_{1}\otimes w_{1}^{\prime})$. Thus $\mu:\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W}\to\mathcal{W}$ satisfies the assumptions of 2.31, so $\mu$ is a $\mathcal{V}$-internal left adjoint, as was to be shown.
Conversely, suppose $\mathcal{W}$ is rigid. In this case, it is dualizable (by definition or by 4.55) and the canonical map $\mathrm{trcl}_{\mathcal{W}}^{\mathcal{V}}\to{\mathrm{at}}_{\mathcal{W}}^{\mathcal{V}}$ is an equivalence by 4.52 and Example 2.9. Therefore 2.25 shows that every object in $\mathcal{W}$ is trace-class presentable. ∎
Construction 4.96.
@@
We define $\mathcal{W}^{(\alpha)}$ by induction as follows. $\mathcal{W}^{(0)}=\mathcal{W}$, $\mathcal{W}^{(\alpha+1)}$ is generated under colimits and tensor products by objects $x\in\mathcal{W}^{(\alpha)}$ that can be written as $\mathrm{colim}_{I}^{W}x_{i}$ with $x_{i}\in(\mathcal{W}^{(\alpha)})^{\kappa}$ and $W\to\mathrm{trcl}^{\mathcal{V}}_{\mathcal{W}^{(\alpha)}}(x_{\bullet},x)$, $W$ a $\kappa$-compact weight.
@@
For the same reason as in 2.43, for $\alpha$ large enough $\mathcal{W}^{(\alpha+1)}=\mathcal{W}^{(\alpha)}$.
@@
This follows at once from 4.94. ∎
Observation 4.98.
@@
The “in particular” part follows from the fact that the functor $\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})\to\mathcal{W}$ induces an equivalence $\mathrm{Rig}_{\mathcal{V}}(\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa}))\xrightarrow{\simeq}\mathrm{Rig}_{\mathcal{V}}(\mathcal{W})$ by 4.67.
@@
Note that by comparing their restriction to $\mathcal{W}^{\kappa}$, we find that $p_{\mathcal{V}}\circ\mathcal{P}_{\mathcal{V}}(\mathrm{trcl}(m,-)):\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})\to\mathcal{V}$ is exactly $\mathrm{trcl}(y(m),-)$ so that $\mathrm{trcl}(y(m),\hat{y}(-))$ is the universal $\mathcal{V}$-linear approximation to $\mathrm{trcl}(m,-)$, by 2.11.
@@ 5 Presentability results
In this section, we prove an analogue of our main theorem for rigid categories, namely we prove that the category of rigid $\mathcal{V}$-algebras is presentable. In the case of $\mathcal{V}=\mathbf{Sp}$, we have an explicit description of the right adjoint $\mathrm{CAlg}^{\mathrm{rig}}_{\mathbf{Sp}}\to\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st},\kappa})$ by 4.77 which allows to give a quick proof of presentability, so we first deal with this special case.
In general, we have to work harder and give a “parametrized” version of [Ram24, Theorem B]. The same strategy also works for a version of this result with locally rigid categories, although it is not as clean because maps between locally rigid categories are not automatically internal left adjoints.
@@ 5.1 Over $\mathbf{Sp}$
Unlike in the case of dualizable categories, we do not have to restrict the morphisms, because any $\mathbf{Sp}$-algebra map between rigid $\mathbf{Sp}$-algebras is an internal left adjoint, cf. 4.48.
@@
(1) We start by noting that, as a full subcategory of $B$, $A$ consists exactly of those $b$’s such that for any $p:x\to y$ such that $i^{R}(p)$ is an equivalence, $\operatorname{Map}(b,x)\to\operatorname{Map}(b,y)$ is also an equivalence.
@@
Proof of 5.1.
By [Ram24, Theorem 3.4] and 4.48, we have an inclusion $\mathrm{CAlg}^{\mathrm{rig}}\subset\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st},\omega_{1}})$, and this furthermore has a right adjoint $\mathrm{Rig}(-)$ by 4.77. By 5.3, it suffices to prove that $\mathrm{Rig}(-)$, viewed as an endofunctor of $\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st},\omega_{1}})$, is accessible.
@@
Finally, let us note that if the unit of $\mathcal{M}$ is $\kappa$-compact, then “$f:x\to y$ is a trace-class map in $\mathcal{M}$” can be witnessed by a $\kappa$-compact object $d\in\mathcal{M}$, in the notation of 3.1. It follows that on $\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st},\omega_{1}})$, the full subfunctor of $\mathcal{M}\mapsto(\mathcal{M}^{\omega_{1}})^{\mathbb{Q}_{\geq 0}}$ spanned by those diagrams where all transition maps are trace-class is an $\omega_{1}$-accessible subfunctor.
@@ 5.2 In general
Similarly, we prove:
Theorem 5.5.
Let $\mathrm{CAlg}^{\mathrm{loc.rig}}_{\mathcal{V}}\subset\mathrm{CAlg}(\mathbf{Mod}_{\mathcal{V}}(\mathbf{Pr}^{\mathrm{L}}_{\kappa}))$ be the non-full subcategory spanned by locally rigid $\mathcal{V}$-algebras and internal left adjoints between them. $\mathrm{CAlg}^{\mathrm{loc.rig}}_{\mathcal{V}}$ is presentable.
Remark 5.6.
We do not know whether the same holds for the category of locally rigid $\mathcal{V}$-algebras and functors that preserve $\kappa$-compacts, for a fixed $\kappa$. The proof from 4.19 seems not to apply anymore, and so we do not even know whether it admits all colimits.
We have seen in 4.19 and its variant 4.53 that these categories admit all small colimits. Thus this really is a claim about accessibility. In both cases, we will witness the relevant categories as pullbacks of accessible categories along accessible functors.
@@
[…]
@@
[…]
@@
We have explained the first part in 5.12 as a consequence of [Ram24, Theorem B].
@@
Proof of 5.4 and 5.5 .
Fix $\kappa\geq\omega_{1}$ so that $\mathcal{V}\in\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\kappa})$. By 4.53 and 4.19, $\mathrm{CAlg}^{\mathrm{rig}}_{\mathcal{V}}$ and $\mathrm{CAlg}_{\mathcal{V}}^{\mathrm{loc.rig}}$ have all colimits, and so the theorem is about accessibility.
Any (locally) rigid $\mathcal{V}$-algebra is dualizable over $\mathcal{V}$ and hence $\kappa$-compactly generated by [Ram24, Theorem 3.4].
@@
where $\mathbf{Mod}^{\mathcal{V}}_{\kappa},\mathbf{Mod}^{\mathcal{V}}_{\kappa,{\mathrm{dbl}}}$ are as in 5.9 and the discussion following it, and the functor $\mathrm{CAlg}(\mathbf{Mod}(\mathcal{V})^{\mathrm{dbl}})\to(\mathbf{Mod}^{\mathcal{V}}_{\kappa})^{\Delta^{1}}$ is the functor $\mathcal{W}\mapsto(\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W}\to\mathcal{W})$.
Now, from the description of colimits in $\mathbf{Mod}^{\mathcal{V}}_{\kappa}$ (cf. the proof of 5.13), one can check that this functor is accessible, and hence 5.13 guarantees that the pullback is also accessible, as was needed.
For $\mathrm{CAlg}_{\mathcal{V}}^{\mathrm{loc.rig}}$, only little modification is needed. Namely, we simply need to point out that $\mathrm{CAlg}(\mathbf{Mod}_{\mathcal{V}}(\mathbf{Pr}^{\mathrm{L}}_{\kappa}))\times_{\mathbf{Mod}_{\mathcal{V}}(\mathbf{Pr}^{\mathrm{L}}_{\kappa})}\mathbf{Mod}(\mathcal{V})^{\mathrm{dbl}}$ is presentable, and we can replace $\mathrm{CAlg}(\mathbf{Mod}(\mathcal{V})^{\mathrm{dbl}})$ in the previous proof by this category, and now the pullback is $\mathrm{CAlg}_{\mathcal{V}}^{\mathrm{loc.rig}}$. ∎
We now point out that 5.4 allows us to prove more abstractly the existence of rigidifications, by combining the following two corollaries (which of course, we now logically do not need):
@@
Since any rigid $\mathcal{V}$-algebra is itself in $\mathrm{CAlg}(\mathbf{Mod}_{\mathcal{V}}(\mathbf{Pr}^{\mathrm{L}}_{\kappa}))$ by [Ram24, Theorem 3.4], this is really about the map $\mathcal{W}_{0}\to\mathcal{W}$ itself preserving $\kappa$-compacts. But note that this map is the composite: $\mathcal{W}_{0}\to\mathcal{W}_{0}\otimes_{\mathcal{V}}\mathcal{W}\to\mathcal{W}$, where the second map is an internal left adjoint by rigidity of $\mathcal{W}_{0}$ and by 4.11, while the first map is $\mathcal{W}_{0}\otimes_{\mathcal{V}}(\mathcal{V}\to\mathcal{W}_{0})$. Since the latter is in $\mathbf{Mod}_{\mathcal{V}}(\mathbf{Pr}^{\mathrm{L}}_{\kappa})$ by assumption, the tensor product is in there as well, and so we are done (this is simply a variation on 4.48). ∎
In particular, the common value of all the $\mathrm{Rig}_{\mathcal{V},\kappa}(\mathcal{W})$ for $\kappa$ large enough is the rigidification of $\mathcal{W}$ in the sense of 4.61.

### v2

@@ 4.3.1 Over $\mathbf{Sp}$
Since $\mathcal{V}_{i}$ is rigid, this implies that it is trace-class, and thus the whole of $(x_{i}^{r})_{i}\to(x_{i}^{s})_{i}$ is trace-class in $\prod_{i}\mathcal{V}_{i}$. By the usual trick, it implies that $\mathrm{colim}_{r\in\mathbb{Q}_{\geq 0}}y((x_{i}^{r})_{i})$ is also a $\mathbb{Q}_{\geq 0}$-indexed colimit of things in $\prod_{i}^{\mathrm{dbl}}\mathcal{V}_{i}$ along trace-class maps, and thus it follows that $\prod_{i}^{\mathrm{dbl}}\mathcal{V}_{i}$ is rigid by Section˜4.2.
Finite limits: The empty limit is the terminal object, i.e. 0, and this is clearly rigid. So we are left with pullbacks: let $\mathcal{V}_{0}\xrightarrow{p_{0}}\mathcal{V}_{01}\xleftarrow{p_{1}}\mathcal{V}_{1}$ be a span in $\mathrm{CAlg}^{\mathrm{rig}}$, and $\mathcal{V}$ be its pullback in $\mathrm{CAlg}((\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})^{\mathrm{dbl}})$, $\mathcal{V}^{\prime}$ its pullback in $\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st}})$, and $\overset{\rightarrow}{\mathcal{V}}$ the doubly lax pullback as in Section˜4.3.1.
@@
Furthermore, we have observed in Section˜4.3.1 that the bottom left vertical map is an isomorphism, and it is clear from Section˜4.2 that the top left vertical map also is. Thus it suffices to prove that the image of $\hat{y}:\mathcal{V}\to\mathrm{Ind}(\mathcal{V})\to\mathrm{Ind}(\overset{\rightarrow}{\mathcal{V}})$ is contained in the image of $\mathrm{Rig}(\overset{\rightarrow}{\mathcal{V}})$.
However we already know that the image of $\mathcal{V}$ in $\mathrm{Ind}(\mathcal{V}^{\prime})$ is generated under colimits by $\mathbb{Q}$-indexed colimits of maps between representables whose projections to each $\mathcal{V}_{i}$ are trace-class. Since each map in $\mathbb{Q}$ is a composite of two maps, it follows from Section˜4.3.1 that each of the maps in question is trace-class in $\overset{\rightarrow}{\mathcal{V}}$, and hence, that the relevant $\mathbb{Q}$-colimits are in $\mathrm{Rig}(\overset{\rightarrow}{\mathcal{V}})$, as was to be shown. ∎
@@ 4.3.2 In general
We will prove in Section˜5 that the category of rigid $\mathcal{V}$-algebras is presentable, and explain how to obtain from this a different proof of the existence of rigidifications. However, it is still not explicit enough to obtain ˜4.67 - the goal of this subsection is to provide an actual construction of the rigidification for which we have enough control to prove this result. This construction is motivated by the construction over $\mathbf{Sp}$, and just as in this case, we first need an “intrinsic” description of rigid categories, obtained as ˜4.98.
[…]
We already noted in Section˜4.3 that if $\mathbf{1}_{\mathcal{W}}$ is atomic, then $\mathrm{Rig}_{\mathcal{V}}(\mathcal{W})\to\mathcal{W}$ is fully faithful, so our goal is to describe the essential image.
@@
Suppose $\mathcal{W}$ is a commutative $\mathcal{V}$-algebra with $\mathcal{V}$-atomic unit. $\mathcal{W}$ is rigid if and only if it satisfies an analogue of ˜2.25 with trace-class presentability, that is, if and only if:
@@
As in the case of ˜2.25, it is worth starting with a preparatory lemma:
@@
Section˜4.3.2 shows that if $f:\mathcal{M}\to\mathcal{N}$ is a lax $\mathcal{W}$-linear functor which is strong $\mathcal{V}$-linear and colimit-preserving, then any of the assumptions 1.-4. or their “generated by” analogues imply that $f$ is strong $\mathcal{W}$-linear.
@@
First we observe that there is a map $\mathrm{trcl}_{\mathcal{W}}^{\mathcal{V}}\to{\mathrm{at}}_{\mathcal{W}}^{\mathcal{V}}$ since $\mathbf{1}_{\mathcal{W}}$ is atomic. Therefore, our assumptions imply the assumptions of ˜2.25, so that $\mathcal{W}$ is dualizable over $\mathcal{V}$. We may therefore apply Section˜2 to check whether this is a $\mathcal{V}$-internal left adjoints.
Now we note that tensor products of trace-class maps are trace-class - a refinement of this is the existence of a transformation $\mathrm{trcl}_{\mathcal{W}}^{\mathcal{V}}(w_{0},w_{1})\otimes\mathrm{trcl}_{\mathcal{W}}^{\mathcal{V}}(w_{0}^{\prime},w_{1}^{\prime})\to\mathrm{trcl}_{\mathcal{W}}^{\mathcal{V}}(w_{0}\otimes w_{0}^{\prime},w_{1}\otimes w_{1}^{\prime})$. Thus $\mu:\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W}\to\mathcal{W}$ satisfies the assumptions of Section˜2, so $\mu$ is a $\mathcal{V}$-internal left adjoint, as was to be shown.
Conversely, suppose $\mathcal{W}$ is rigid. In this case, it is dualizable (by definition or by Section˜4.2) and the canonical map $\mathrm{trcl}_{\mathcal{W}}^{\mathcal{V}}\to{\mathrm{at}}_{\mathcal{W}}^{\mathcal{V}}$ is an equivalence by Section˜4.2 and Section˜2. Therefore ˜2.25 shows that every object in $\mathcal{W}$ is trace-class presentable. ∎
Construction 4.100.
@@
We define $\mathcal{W}^{(\alpha)}$ by induction as follows. $\mathcal{W}^{(0)}=\mathcal{W}$, $\mathcal{W}^{(\alpha+1)}$ is generated under colimits and tensor products by objects $x\in\mathcal{W}^{(\alpha)}$ that can be written as $\mathrm{colim}_{I}^{W}x_{i}$ with $x_{i}\in(\mathcal{W}^{(\alpha)})^{\kappa}$ and $\to\mathrm{trcl}^{\mathcal{V}}_{\mathcal{W}^{(\alpha)}}(x_{\bullet},x), WW$ a $\kappa$-compact weight.
@@
For the same reason as in Section˜2, for $\alpha$ large enough $\mathcal{W}^{(\alpha+1)}=\mathcal{W}^{(\alpha)}$.
@@
This follows at once from ˜4.98. ∎
Observation 4.102.
@@
The “in particular” part follows from the fact that the functor $\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa})\to\mathcal{W}$ induces an equivalence $\mathrm{Rig}_{\mathcal{V}}(\mathcal{P}_{\mathcal{V}}(\mathcal{W}^{\kappa}))\xrightarrow{\simeq}\mathrm{Rig}_{\mathcal{V}}(\mathcal{W})$ by Section˜4.3.
@@
[…]
@@ 5 Presentability results
In this section, we prove an analogue of our main theorem for rigid categories, namely we prove that the category of rigid $\mathcal{V}$-algebras is presentable. In the case of $\mathcal{V}=\mathbf{Sp}$, we have an explicit description of the right adjoint $\mathrm{CAlg}^{\mathrm{rig}}_{\mathbf{Sp}}\to\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st},\kappa})$ by ˜4.81 which allows to give a quick proof of presentability, so we first deal with this special case.
In general, we have to work harder and give a “parametrized” version of [RAM24, Theorem B].
@@ 5.1 Over $\mathbf{Sp}$
Unlike in the case of dualizable categories, we do not have to restrict the morphisms, because any $\mathbf{Sp}$-algebra map between rigid $\mathbf{Sp}$-algebras is an internal left adjoint, cf. Section˜4.2.
@@
(1) We start by noting that, as a full subcategory of $B$, AA consists exactly of those $b$’s such that for any $p:x\to y$ such that $i^{R}(p)$ is an equivalence, $\operatorname{Map}(b,x)\to\operatorname{Map}(b,y)$ is also an equivalence.
@@
Proof of ˜5.1.
By [RAM24, Theorem 3.4] and Section˜4.2, we have an inclusion $\mathrm{CAlg}^{\mathrm{rig}}\subset\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st},\omega_{1}})$, and this furthermore has a right adjoint $\mathrm{Rig}(-)$ by ˜4.81. By Section˜5.1, it suffices to prove that $\mathrm{Rig}(-)$, viewed as an endofunctor of $\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st},\omega_{1}})$, is accessible.
@@
Finally, let us note that if the unit of $\mathcal{M}$ is $\kappa$-compact, then $f:x\to y$ is a trace-class map in $\mathcal{M}$” can be witnessed by a $\kappa$-compact object $d\in\mathcal{M}$, in the notation of Section˜3. It follows that on $\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\mathrm{st},\omega_{1}})$, the full subfunctor of $\mathcal{M}\mapsto(\mathcal{M}^{\omega_{1}})^{\mathbb{Q}_{\geq 0}}$ spanned by those diagrams where all transition maps are trace-class is an $\omega_{1}$-accessible subfunctor.
@@ 5.2 In general
Warning 5.5.
An earlier version of this paper contained a similar claim for a variant of the category of locally rigid $\mathcal{V}$-algebras. The claim in question was mistaken, as was pointed out to the author by Jiacheng Liang, and currently we do not know a way to make any such claim both nontrivial and true. Such a result might be provable if one could prove an analogue of Section˜4.1 where the functors are only required to preserve $\kappa$-compacts for some fixed $\kappa$.
We have seen in Section˜4.1 that $\mathrm{CAlg}^{\mathrm{rig}}_{\mathcal{V}}$ admits all small colimits. Thus this really is a claim about accessibility. We will prove this by witnessing the category of rigid commutative algebras as a pullback of accessible categories along accessible functors.
@@
Let $B$ be a presentable category, and $e:B\to\mathbf{Pr}^{\mathrm{L}}$ be an accessible functor. The unstraightened cocartesian fibration $E\xrightarrow{p}B$ has a presentable total category $E$.
@@
Let $\mathbf{Mod}^{\mathcal{V}}_{\kappa}\xrightarrow{p}\mathrm{CAlg}(\mathbf{Mod}_{\mathcal{V}}(\mathbf{Pr}^{\mathrm{L}}_{\kappa}))$ be the cocartesian fibration classifying the functor $\mathcal{W}\mapsto\mathbf{Mod}_{\mathcal{W}}(\mathbf{Pr}^{\mathrm{L}}_{\kappa})$.
@@
We have explained the first part in Section˜5.2 as a consequence of [RAM24, Theorem B].
@@
Proof of ˜5.4.
Fix $\kappa\geq\omega_{1}$ so that $\mathcal{V}\in\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}}_{\kappa})$. By Section˜4.2, $\mathrm{CAlg}^{\mathrm{rig}}_{\mathcal{V}}$ has all colimits, and so the theorem is about accessibility.
Any rigid $\mathcal{V}$-algebra is dualizable over $\mathcal{V}$ and hence $\kappa$-compactly generated by [RAM24, Theorem 3.4].
@@
where $\mathbf{Mod}^{\mathcal{V}}_{\kappa},\mathbf{Mod}^{\mathcal{V}}_{\kappa,{\mathrm{dbl}}}$ are as in Section˜5.2 and the discussion following it, and the functor $\mathrm{CAlg}(\mathbf{Mod}(\mathcal{V})^{\mathrm{dbl}})\to(\mathbf{Mod}^{\mathcal{V}}_{\kappa})^{\Delta^{1}}$ is the functor $\mathcal{W}\mapsto(\mathcal{W}\otimes_{\mathcal{V}}\mathcal{W}\to\mathcal{W})$.
Now, from the description of colimits in $\mathbf{Mod}^{\mathcal{V}}_{\kappa}$ (cf. the proof of Section˜5.2), one can check that this functor is accessible, and hence Section˜5.2 guarantees that the pullback is also accessible, as was needed. ∎
We now point out that ˜5.4 allows us to prove more abstractly the existence of rigidifications, by combining the following two corollaries (which of course, we now logically do not need):
@@
Since any rigid $\mathcal{V}$-algebra is itself in $\mathrm{CAlg}(\mathbf{Mod}_{\mathcal{V}}(\mathbf{Pr}^{\mathrm{L}}_{\kappa}))$ by [RAM24, Theorem 3.4], this is really about the map $\mathcal{W}_{0}\to\mathcal{W}$ itself preserving $\kappa$-compacts. But note that this map is the composite: $\mathcal{W}_{0}\to\mathcal{W}_{0}\otimes_{\mathcal{V}}\mathcal{W}\to\mathcal{W}$, where the second map is an internal left adjoint by rigidity of $\mathcal{W}_{0}$ and by Section˜4.1, while the first map is $\mathcal{W}_{0}\otimes_{\mathcal{V}}(\mathcal{V}\to\mathcal{W}_{0})$. Since the latter is in $\mathbf{Mod}_{\mathcal{V}}(\mathbf{Pr}^{\mathrm{L}}_{\kappa})$ by assumption, the tensor product is in there as well, and so we are done (this is simply a variation on Section˜4.2). ∎
In particular, the common value of all the $\mathrm{Rig}_{\mathcal{V},\kappa}(\mathcal{W})$ for $\kappa$ large enough is the rigidification of $\mathcal{W}$ in the sense of Section˜4.3.